A company produces a product which has a standard variable production cost of $8 per unit made up as

follows:

                                                               $ Per Unit

Direct material                                 $4.60 (2kg X $2.30 per kg)

Direct labour                                     $2.10 (0.7 hours x $3.00 per hour)

Variable overhead                          $1.3

Fixed manufacturing costs are treated as period costs. The following information is available for the period just ended,

Variable manufacturing cost of sales (at standard cost)                                  $263,520

Opening stock of finished goods (at standard cost)                                         $120,800

Closing stock of finished goods (at standard cost)                                          $146,080

Direct material price variance                                                                          $2,571 U

Raw materials used in manufacture (at actual cost)                                        $170,310

Direct labour rate variance                                                                               $4,760 U

Direct labour efficiency variance                                                                      $3,240 F

Required:

(a)   Determine for the period ended.

(i)              The total actual direct labour cost, and

(ii)              The actual cost per kg of raw material used.

(b)   Outline the possible causes of the raw materials variances
Answer #1
Actual direct labor cost
Variable manufacturing cost of sales $263,520
Add: Ending stock of finished goods (at standard cost)   146080
Less:Opening stock of finished goods (at standard cost)   -120800
cost of goods manufactured $288,800
cost of goods manufactured $288,800
No of units produced*(4.6+2.1+1.3)=288800
Std No of units produced=288800/(4.6+2.1+1.3) 36100
Standard cost of direct labor
36100*2.1 75810
Add: Direct labour rate variance                                                                       4760
Less: Direct labour efficiency variance         -3240
Actual direct labor cost 77330 answer 1
ans 2
Actual cost per kg of raw material used.
Direct material price variance   $2,571
(AQ*AR)-(AQ*Sr)
AQ*SR=170310-2571 167739
Aq*2.3=167739
AQ=167739/2.3 72930
72930*AR=170310
AR=170310/72930 2.34
Actual rate=$2.34
ans 3
Causes of raw material variance;
1) Actual cost of raw material is more or less than budgeted. So when the market price of the material increases or decreases than there could be unfavorable or favorable variance respectively
2) actual quantity used can be less or more than standard quantity allowed. It may be due to inefficient labor in case of unfavorable variance and more skilled labor in case of favorable variance
3) Substandard quality used in production than was planned than there is unfavorable variance
If any doubt please comment